OpenStudy (anonymous):

What is the simplified form of each expression? -(4)^-2

OpenStudy (diyadiya):

\[\Large x^{-m}= \frac{1}{x^m}\]

OpenStudy (diyadiya):

\[\Large -(4)^-2 = \frac{1}{-(4)^2}\]

OpenStudy (diyadiya):

is that clear?

OpenStudy (diyadiya):

so what will be the answer?

OpenStudy (diyadiya):

@Jen1986

OpenStudy (anonymous):

idk how to make it set up like that.... so the answer is 1/-(4)^2?

OpenStudy (anonymous):

I mean i don't know how to make it look like a fraction like you did above

OpenStudy (diyadiya):

Its ok ..but what is -(4)^2?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

?

OpenStudy (diyadiya):

What happend to the negetive ? so -16 so the final answer is?

OpenStudy (diyadiya):

?

OpenStudy (anonymous):

does the negative stay?

OpenStudy (anonymous):

ugh can't type

OpenStudy (anonymous):

I have no idea how to do any of this... this is why I posted the question

OpenStudy (diyadiya):

Yes if it was (-4)^2 then the negetive wouldn't but since it is -(4)^2 the negative stays so 1/-(4)^2 = ?

OpenStudy (anonymous):

-16?

OpenStudy (diyadiya):

1/-16 or -1/16

OpenStudy (anonymous):

Ok thank you

OpenStudy (diyadiya):

no problem :)